OpenStudy (anonymous):

Jeanine bought a new pair of hiking boots on sale for 25% off the regular price. The following formula describes this scenario: s = r - 0.25r Where r = Regular price s = Sale price If the regular price was $216.67, how much did Jeanine pay for the hiking boots?

OpenStudy (anonymous):

@amistre64 Help please?

OpenStudy (anonymous):

@hartnn Help? :L

OpenStudy (texaschic101):

just sub in 216.67 in for r in the formula s = r - 0.25r

OpenStudy (anonymous):

I did, however I get a huge decimal.

OpenStudy (anonymous):

I do this this first, correct? s = r - 0.25 * 216.67 0.25 * 216.67 = 54.1675

OpenStudy (texaschic101):

s = 216.67 - 0.25(216.67) s = 216.67 - 54.1675 rounds to 54.17 s = 216.67 - 54.17 = 162.50

OpenStudy (anonymous):

Okay so I was on the right track :P thanks! :D

OpenStudy (texaschic101):

yes you were :)

OpenStudy (anonymous):

Is it okay if I tag you in the others I have? :)

OpenStudy (texaschic101):

yes...I will help if I can :)

OpenStudy (anonymous):

Okay, thanks again :)

OpenStudy (texaschic101):

no problem :)